# Hybrid RAG-empowered Multi-modal LLM for Secure Healthcare Data Management: A Diffusion-based Contract Theory Approach

### v1

@@ Abstract
Secure data management and effective data sharing have become paramount in the rapidly evolving healthcare landscape. The advancement of generative artificial intelligence has positioned Multi-modal Large Language Models (MLLMs) as crucial tools for managing healthcare data. MLLMs can support multi-modal inputs and generate diverse types of content by leveraging large-scale training on vast amounts of multi-modal data. However, critical challenges persist in developing medical MLLMs, including healthcare data security and freshness issues, affecting the output quality of MLLMs. In this paper, we propose a hybrid Retrieval-Augmented Generation (RAG)-empowered medical MLLMs framework for healthcare data management. This framework leverages a hierarchical cross-chain architecture to facilitate secure data training. Moreover, it enhances the output quality of MLLMs through hybrid RAG, which employs multi-modal metrics to filter various unimodal RAG results and incorporates these retrieval results as additional inputs to MLLMs. Additionally, we employ age of information to indirectly evaluate the data freshness impact of MLLMs and utilize contract theory to incentivize healthcare data holders to share fresh data, mitigating information asymmetry in data sharing. Finally, we utilize a generative diffusion model-based reinforcement learning algorithm to identify the optimal contract for efficient data sharing. Numerical results demonstrate the effectiveness of the proposed schemes, which achieve secure and efficient healthcare data management.
@@ I Introduction
As the backbone of society and human interaction, the healthcare industry has undergone a seismic shift with the advent of digital technology[1]. This transformation has not only revolutionized patient care but also catalyzed the generation, storage, and analysis of vast amounts of healthcare data[2], which encompasses different types of big healthcare data[3], such as omics data, clinical data, electronic health records, sensing data, etc. Although the exponential growth in healthcare data volume holds the potential to revolutionize healthcare by providing insights into patient care, disease patterns, and treatment effectiveness, it also requires sophisticated tools for its analysis and interpretation. Fortunately, Generative Artificial Intelligence (GAI) as a new branch of AI has emerged as a potent technology that can effectively analyze vast datasets and generate various content[4, 5]. Especially, by analyzing patient data, including medical history and treatment outcomes, GAI can contribute to personalized medicine by considering individual patient characteristics and generating customized treatment plans[5, 6].
@@
Large Language Models (LLMs), as a technological application of GAI, can achieve general-purpose language generation and conventional natural language processing tasks, which hold the potential to significantly transform healthcare management. With the integration of multi-modal data into LLMs, patients can effectively comprehend many aspects of their physical health through Multi-modal Large Language Models (MLLMs)[7]. For example, the latest GPT-4, equipped with vision capabilities and exceptional performance in natural language processing tasks, can be fine-tuned as a powerful guidance tool in the healthcare domain[8]. However, the sheer size and complexity of MLLMs necessitate efficient retrieval mechanisms to enhance their performance further. Retrieval-Augmented Generation (RAG) is a cutting-edge technique that boosts the reliability and accuracy of GAI models by retrieving facts from an external knowledge base[9]. Furthermore, RAG can capitalize on the similarity between the alignment vectors of the query to retrieve pertinent data, thereby enhancing user prompts by integrating relevantly retrieved data within the context, enabling MLLMs to generate accurate and contextually appropriate responses[4]. Thanks to the prominent capabilities of RAG, the integration of MLLMs and RAG has been widely used in various domains[10, 11].
@@
Despite the advancements in RAG-empowered MLLMs, there are several persistent challenges in the application of these technologies for healthcare data management: 1) Since healthcare data is normally multi-modal and stored in different databases in a distributed manner, unimodal RAG using a single search manner, such as vector similarity search and keyword search[11], may not efficiently retrieve multi-modal healthcare data to support LLM tasks that handle multiple modes. 2) The application of MLLMs in analyzing healthcare data poses significant security risks and privacy concerns[12]. Healthcare data is highly sensitive, and any breach or misuse can have severe consequences for patients and healthcare providers[2]. Thus, ensuring the confidentiality and integrity of healthcare data during MLLM processing is a critical concern. 3) Pre-trained medical MLLMs can result in inaccurate inferences during task-specific fine-tuning due to biases in the dataset. Hence, incorporating fresh high-quality healthcare data is crucial for fine-tuning MLLMs to avoid incorrect learning patterns[4]. 4) Considering the problem of information asymmetry, healthcare data holders often have more data information, and appropriate incentive mechanisms need to be implemented to encourage healthcare data holders to provide accurate and up-to-date information, which is helpful to enhance the medical diagnostic quality of MLLMs empowered by RAG.
@@
To address these challenges, we propose a hybrid RAG-empowered medical MLLMs framework for healthcare data management. Specifically, we utilize cross-chain techniques to allow participants to share data without the involvement of a central institution[13]. To enhance the diagnostic quality of MLLMs, we leverage hybrid multi-modal RAG to further refine the retrieval results. Compared to RAG-empowered LLMs, we employ multi-modal metrics to filter multiple unimodal RAG results and incorporate these retrieval results into MLLMs as additional inputs. Furthermore, we apply Age of Information (AoI) to indirectly evaluate the quality of healthcare data and propose a contract theory model to encourage participants to share fresh data, thus coping with the information asymmetry of data sharing. Besides, considering the dynamic environment of data sharing, we use Generative Diffusion Model (GDM)-based reinforcement learning to find the optimal contract[14]. The key contributions of this paper are summarized as follows:
@@
  • We develop a novel hybrid RAG-empowered MLLMs framework for healthcare data management. This framework facilitates secure interactions between healthcare data holders and the MLLM service provider using a cross-chain system for secure healthcare data transmission. Through employing hybrid RAG, MLLMs improve their quality and complete specific tasks by employing hybrid RAG to retrieve multi-modal healthcare data.
@@
  • To optimize time-sensitive learning tasks within MLLM services, we apply AoI as a data freshness metric for indirectly assessing the quality of healthcare data. Furthermore, we propose a contract theory model to incentivize healthcare data holders to contribute high-quality healthcare data with small AoI, thus improving the inference performance of hybrid RAG-empowered MLLMs.
@@
  • To tackle the high-dimensional complexity of the formulated problem, we utilize the GDM-based reinforcement learning algorithm to find the optimal contract for efficient data sharing. Numerical results demonstrate the effectiveness of the proposed GDM-based scheme and its superiority over the Deep Reinforcement Learning (DRL)-based scheme.
@@
The remainder of this paper is organized as follows. Section II reviews the related work. In Section III, we propose a hybrid RAG-empowered medical MLLM framework based on cross-chain technology to enhance data management in the healthcare industry. In Section IV, we introduce a contract theory model to motivate healthcare data holders to provide high-quality healthcare data. In Section V, we present a GDM-based algorithm for optimal contract design. Section VI provides a performance analysis of the proposed schemes. Finally, Section VII concludes this paper. The main notations in our article are summarized in Table I.
@@ II Related Work
II-A LLMs for Data Management
Recent advancements in LLMs have significantly contributed to data management, with numerous research efforts focusing on various aspects such as data analysis, predictive modeling, and decision support systems. Some studies use the strong interpretative abilities of LLMs as agents to continuously improve data storage, data analysis, and additional areas[8, 15]. For instance, the authors in [8] introduced GPT-4, which has demonstrated remarkable capabilities in understanding and generating human-like text, facilitating various data management tasks[8]. The authors in [16] presented an LLM-based database framework that leverages LLMs for automatic prompt generation and model fine-tuning, which performs highly effective in query rewriting and index tuning[16]. In [17], the authors introduced Data-Copilot, which is a data analysis agent capable of autonomously querying, processing, and visualizing vast amounts of data to meet various human needs. Existing works mainly focus on unimodal LLMs processing text data, while relatively insufficient research has been done on the integration of multi-modal data such as text, images, and structured data. Addressing this gap could substantially boost the functionality of data management systems, especially in complex and data-intensive fields like healthcare.
II-B RAG-empowered LLMs
@@
RAG has incredible capabilities in enhancing the accuracy and reliability of LLM output by incorporating additional information sources, such as external knowledge bases, and augmenting user prompts with relevant retrieval data in context[9]. As a novel technique, RAG allows LLMs to bypass retraining, allowing access to the most up-to-date information to generate reliable output through retrieval-based generation[10]. In [9], the authors introduced RAG, demonstrating its ability to improve the accuracy and relevance of generated text by incorporating retrieved documents into the generation process. The authors in [18] proposed a hybrid RAG method, which integrates Sentence-Window and Parent-Child approaches, and demonstrated that the proposed method outperforms current state-of-the-art RAG techniques. The authors in [4] introduced a carbon emission optimization framework that integrates RAG and LLMs, making a significant impact on GAI efforts to reduce carbon emissions. Moreover, RAG is gradually emerging as a promising tool for healthcare applications, for example, it can optimize the interpretation of clinical guidelines for liver disease with the help of external medical knowledge[19]. In addition, the authors in [20] retrieved similar image-text pairs based on image-text contrast similarity, and utilized the retrieval attention module to blend the representation of images and questions with the retrieved images and texts, demonstrating effectiveness in simple biomedical visual question answering.
@@ II-C Contract Theory for Data Sharing
Contract theory is a branch of economics that studies how contractual arrangements can be designed to align incentives between parties with asymmetric information[21], and it has been widely used in wireless communication, AI, and other fields[21, 13]. In the context of data sharing, information asymmetry often arises because data holders possess more information about the data than data users. Contract theory can effectively incentivize data sharing by ensuring that both parties benefit from the exchange[13]. For example, the authors in [22] proposed a two-period incentive mechanism for healthcare applications, which takes into account the Willingness To Participate (WTP) of users and satisfies intertemporal incentive compatibility. This dynamic contract design meets essential constraints and achieves higher profits compared to a uniform pricing scheme. In the context of a mobile AI-generated content network with Unmanned Aerial Vehicles (UAVs), the authors in [23] proposed an AoI-based contract theory model to incentivize the contribution of fresh data between UAVs. The authors in [24] proposed an effective incentive mechanism. This mechanism integrates reputation and contract theory to motivate high-reputation mobile devices with high-quality data to engage in model learning in a federated learning scenario. In addition to the above work, several efforts have been made to develop contract theory models under prospect theory to facilitate user-centric sensing data sharing[13]. Despite these contributions, the application of diffusion-based contract theory to incentivize data sharing remains unexplored.
@@ III Hybrid RAG-empowered Medical MLLM Framework
In this section, we propose a hybrid RAG-empowered medical MLLM framework. The detailed methodologies for cross-chain interaction in MLLM training and the utilization of hybrid RAG-empowered MLLM agents for data management are discussed in the following subsections.
@@ III-A Cross-Chain Interaction in MLLM Training
In the health center, robust aggregate MLLMs are developed by training on vast amounts of high-quality multi-modal healthcare data[20]. During the data collection, the health center collects healthcare data sourced from hospitals in diverse regions. However, considering privacy concerns, hospitals may be reluctant to upload all healthcare data to the central health center, influenced by patient willingness and reward structures[13]. Due to the significant computing power required for training MLLMs, the absence of high-performance computing equipment at most hospitals hinders the feasibility of applying federated learning to this framework[25], necessitating alternative approaches to data management and model training. To this end, we utilize cross-chain technology to allow hospitals to securely upload healthcare data and make transactions with the health center.
@@
As shown in Fig. 1, a robust health center utilizes a main chain to manage the comprehensive collection of healthcare data and model updates, and multiple subchains are employed to handle specific tasks from hospitals in diverse regions. These subchains leverage the Internet of Medical Things to gather real-time healthcare data from patients and also manage rewards, model configurations, and additional tasks within the workflow[4]. The main chain ensures centralized oversight, while the sub-chains enable efficient and secure data management operations. Specifically, the main chain $M$ sends data collection tasks to the relay chain $R$, and when the subchain $S$ receives the task, hospitals will upload local multi-modal healthcare data based on the selected contract (Step a). Upon successful verification of cross-chain requests by the miners of the relay chain $R$, the relay chain $R$ returns a readiness confirmation, allowing subchains to upload the multi-modal healthcare data (Step b) [13]. After transmitting the multi-modal healthcare data to the main chain $M$, the health center initiates the training of the MLLM (Step c). Once MLLM training is completed, the health center can access the MLLM API through the relay chain $R$ (Step d), and the health center rewards them with monetary compensation based on their data contribution (Step e)[13].
@@
Figure 1: The overview of the hybrid RAG-empowered medical MLLM framework for healthcare data management. Part A shows the cross-chain interaction for secure healthcare data sharing. Part B depicts the processes of multi-modal input optimization based on a hybrid multi-modal RAG module. Part C presents the framework of MLLM inference based on the multi-modal healthcare data.
III-B Hybrid RAG-empowered MLLMs Agents for Data Management

### v2

# Hybrid RAG-empowered Multi-modal LLM for Secure Data Management in Internet of Medical Things: A Diffusion-based Contract Approach

@@ Abstract
Secure data management and effective data sharing have become paramount in the rapidly evolving healthcare landscape, especially with the growing integration of the Internet of Medical Things (IoMT). The rise of generative artificial intelligence has further elevated Multi-modal Large Language Models (MLLMs) as essential tools for managing and optimizing healthcare data in IoMT. MLLMs can support multi-modal inputs and generate diverse types of content by leveraging large-scale training on vast amounts of multi-modal data. However, critical challenges persist in developing medical MLLMs, including security and freshness issues of healthcare data, affecting the output quality of MLLMs. To this end, in this paper, we propose a hybrid Retrieval-Augmented Generation (RAG)-empowered medical MLLM framework for healthcare data management. This framework leverages a hierarchical cross-chain architecture to facilitate secure data training. Moreover, it enhances the output quality of MLLMs through hybrid RAG, which employs multi-modal metrics to filter various unimodal RAG results and incorporates these retrieval results as additional inputs to MLLMs. Additionally, we employ age of information to indirectly evaluate the data freshness impact of MLLMs and utilize contract theory to incentivize healthcare data holders to share their fresh data, mitigating information asymmetry during data sharing. Finally, we utilize a generative diffusion model-based deep reinforcement learning algorithm to identify the optimal contract for efficient data sharing. Numerical results demonstrate the effectiveness of the proposed schemes, which achieve secure and efficient healthcare data management.
@@ I Introduction
The healthcare system has seen rapid advancements with the integration of advanced technologies like cloud computing, the Internet of Things (IoT), and Artificial Intelligence (AI). These innovations have transformed the sector, giving rise to the Internet of Medical Things (IoMT), which is an interconnected network of medical devices and applications that collect and transmit vital healthcare data[1]. IoMT has not only paved the way for more intelligent and efficient healthcare systems, but also catalyzed the generation, storage, and analysis of vast amounts of big healthcare data[2], including omics data, clinical records, electronic health records, etc[3]. Although the exponential growth in healthcare data volume holds the potential to revolutionize the healthcare industry by providing insights into patient care, disease patterns, and treatment effectiveness, it also requires sophisticated tools for its analysis and interpretation. Fortunately, Generative Artificial Intelligence (GenAI) as a new branch of AI has emerged as a potent technology within IoT landscape[4, 5], enabling the effective analysis of vast datasets and generation of diverse content[6, 7]. In particular, GenAI enhances data management by analyzing complex patient records and treatment data, enabling more efficient sharing of critical healthcare information[8].
@@
Large Language Models (LLMs), as a technological application of GenAI, can achieve general-purpose language generation and conventional natural language processing tasks, which hold the potential to significantly transform healthcare data management in IoMT[9]. With the integration of multi-modal data into LLMs, patients can effectively comprehend many aspects of their physical health through Multi-modal LLMs (MLLMs)[10]. For example, the latest GPT-4, equipped with vision capabilities and exceptional performance in natural language processing tasks, can be fine-tuned as a powerful guidance tool in the healthcare domain[11]. However, the sheer size and complexity of MLLMs necessitate efficient retrieval mechanisms to enhance their performance further. Retrieval-Augmented Generation (RAG) is a cutting-edge technique that boosts the reliability and accuracy of GenAI models by retrieving facts from an external knowledge base[12]. Furthermore, RAG can capitalize on the similarity between the alignment vectors of the query to retrieve pertinent data, thereby enhancing user prompts by integrating relevantly retrieved data within the context, enabling MLLMs to generate accurate and contextually appropriate responses[6]. Thanks to the prominent capabilities of RAG, the integration of MLLMs and RAG has been widely used in various domains[13, 14].
@@
Despite the advancements in RAG-empowered MLLMs, there are several persistent challenges in the application of these technologies for healthcare data management in IoMT: 1) Since healthcare data is normally multi-modal and stored in different databases in a distributed manner, unimodal RAG using a single search manner, such as vector similarity search and keyword search[14], may not efficiently retrieve multi-modal healthcare data to support LLM tasks that handle multiple modes. 2) The application of MLLMs in analyzing healthcare data poses significant security risks and privacy concerns[15]. Healthcare data is highly sensitive, and any breach or misuse can have severe consequences for patients and healthcare providers[2]. Thus, ensuring the confidentiality and integrity of healthcare data during MLLM processing is a critical concern. 3) Pre-trained medical MLLMs can result in inaccurate inferences during task-specific fine-tuning due to biases in the dataset. Hence, incorporating fresh high-quality healthcare data is crucial for fine-tuning MLLMs to avoid incorrect learning patterns[6]. 4) Considering the problem of information asymmetry, healthcare data holders often have more data information, and appropriate incentive mechanisms need to be implemented to encourage healthcare data holders to provide accurate and up-to-date information, which is helpful to enhance the medical diagnostic quality of MLLMs empowered by RAG.
@@
To address these challenges, we propose a hybrid RAG-empowered medical MLLM framework for healthcare data management in IoMT. Specifically, we allow participants to share data without the involvement of a central institution by implementing cross-chain techniques, which support secure and efficient data or asset transfers across multiple chains, effectively mitigating single-point-of-failure risks and enhancing overall security[16]. To enhance the diagnostic quality of MLLMs, we leverage hybrid multi-modal RAG to further refine the retrieval results. Compared with RAG-empowered LLMs, we employ multi-modal metrics to filter multiple unimodal RAG results and incorporate these retrieval results into MLLMs as additional inputs. Furthermore, we apply Age of Information (AoI) to indirectly evaluate the quality of healthcare data and utilize a contract theory model to encourage participants to share fresh data, thus coping with the information asymmetry of data sharing. Besides, considering the dynamic environment of data sharing, we use Generative Diffusion Model (GDM)-based Deep Reinforcement Learning (DRL) algorithms to efficiently find the optimal contract[17]. The key contributions of this paper are summarized as follows:
@@
  • We develop a novel hybrid RAG-empowered MLLM framework for healthcare data management in IoMT. This framework facilitates secure interactions between healthcare data holders and the MLLM service provider using a cross-chain system for secure healthcare data transmission, and MLLMs can improve their quality and complete specific tasks by employing hybrid RAG to retrieve multi-modal healthcare data.
@@
  • To optimize time-sensitive learning tasks within MLLM services, we apply AoI as a data freshness metric to indirectly assess the quality of healthcare data. Furthermore, we formulate a contract theory model to incentivize healthcare data holders to contribute high-quality healthcare data with small AoI, thus improving the inference performance of hybrid RAG-empowered MLLMs.
@@
  • To tackle the high-dimensional complexity of the formulated problem, we employ GDM-based DRL algorithms to determine the optimal contract for efficient data sharing. Numerical results demonstrate the effectiveness of the proposed GDM-based scheme, showing a $20.35\%$ performance improvement over DRL-based schemes, highlighting its superiority in this paper.
@@
The remainder of this paper is organized as follows. Section II reviews the related work. In Section III, we propose a hybrid RAG-empowered medical MLLM framework based on cross-chain technology to enhance data management in IoMT. In Section IV, we introduce a contract theory model to motivate healthcare data holders to provide high-quality healthcare data. In Section V, we present GDM-based DRL algorithms for optimal contract design. Section VI provides a performance analysis of the proposed schemes. Finally, Section VII concludes this paper. The main notations in our article are summarized in Table I.
@@ II Related Work
II-A RAG-empowered LLMs
@@
RAG has incredible capabilities in enhancing the accuracy and reliability of LLM output by incorporating additional information sources, such as external knowledge bases, and augmenting user prompts with relevant retrieval data in context[12]. As a novel technique, RAG allows LLMs to bypass retraining, allowing access to the most up-to-date information to generate reliable output through retrieval-based generation[13]. In [12], the authors introduced RAG, demonstrating its ability to improve the accuracy and relevance of generated text by incorporating retrieved documents into the generation process. The authors in [18] proposed a hybrid RAG method that integrates Sentence-Window and Parent-Child approaches and demonstrated that the proposed method outperforms current state-of-the-art RAG techniques. The authors in [6] introduced a carbon emission optimization framework that integrates RAG and LLMs, making a significant impact on GenAI efforts to reduce carbon emissions. Moreover, RAG is gradually emerging as a promising tool for healthcare applications, for example, it can optimize the interpretation of clinical guidelines for liver disease with the help of external medical knowledge[19]. In addition, the authors in [20] retrieved similar image-text pairs based on image-text contrast similarity, and utilized the retrieval attention module to blend the representation of images and questions with the retrieved images and texts, demonstrating effectiveness in simple biomedical visual question answering.
II-B LLMs for Data Management
IoMT has significantly improved healthcare data management by enabling the seamless collection, transmission, and analysis of vast amounts of patient data through interconnected devices[8]. LLMs further enhance this process by offering advanced capabilities in processing unstructured data, extracting valuable insights, and supporting decision-making with greater efficiency and accuracy[9]. Recent advancements in LLMs have significantly contributed to data management, with numerous research efforts focusing on various aspects such as data analysis, predictive modeling, and decision support systems. Some studies use the strong interpretative abilities of LLMs as agents to continuously improve data storage, data analysis, and additional areas[11, 21]. For instance, the authors in [11] introduced GPT-4, which has demonstrated remarkable capabilities in understanding and generating human-like text, facilitating various data management tasks[11]. The authors in [22] presented an LLM-based database framework that leverages LLMs for automatic prompt generation and model fine-tuning, which performs highly effective in query rewriting and index tuning[22]. In [23], the authors introduced Data-Copilot, which is a data analysis agent capable of autonomously querying, processing, and visualizing vast amounts of data to meet various human needs. Existing works mainly focus on unimodal LLMs processing text data, while relatively insufficient research has been done on the integration of multi-modal data such as text, images, and structured data. Addressing this gap could substantially boost the functionality of data management systems, especially in complex and data-intensive IoMT.
@@ II-C Contract Theory for Data Sharing
Contract theory is a branch of economics that studies how contractual arrangements can be designed to align incentives between parties with asymmetric information[24], and it has been widely used in wireless communication, AI, and other fields[24, 16]. In the context of data sharing, information asymmetry often arises because data holders possess more information about the data than data users. Contract theory can effectively incentivize data sharing by ensuring that both parties benefit from the exchange[16]. For example, the authors in [25] proposed a two-period incentive mechanism for healthcare applications, which takes into account the Willingness To Participate (WTP) of users and satisfies intertemporal incentive compatibility. This dynamic contract design meets essential constraints and achieves higher profits compared to a uniform pricing scheme. In the context of a mobile AI-generated content network with Unmanned Aerial Vehicles (UAVs), the authors in [26] proposed an AoI-based contract theory model to incentivize the contribution of fresh data between UAVs. The authors in [27] proposed an effective incentive mechanism. This mechanism integrates reputation and contract theory to motivate high-reputation mobile devices with high-quality data to engage in model learning in a federated learning scenario. In addition to the above work, several efforts have been made to develop contract theory models under prospect theory to facilitate user-centric sensing data sharing[16].
Despite these advancements, the application of diffusion-based contract theory for data sharing remains unexplored. While existing contract theory models are valuable, they often fall short of addressing the intricate challenges associated with the dynamic and multifaceted nature of data sharing. In many real-world contexts, data sharing is not a simple transaction but a process spanning multiple stages and involving diverse participants[28]. Diffusion-based contract theory incorporates the spread and evolution of information over time and space, enabling a deeper understanding of incentives and behaviors among data holders and users[17]. By accounting for diffusion patterns, this approach allows for more accurate assessments of uncertainties and risks tied to data sharing. Consequently, diffusion-based contract models enable the design of adaptive and effective contractual arrangements that support efficient and sustainable data sharing.
@@ III Hybrid RAG-empowered Medical MLLM Framework
In this section, we propose a hybrid RAG-empowered medical MLLM framework in IoMT. The detailed methodologies for cross-chain interaction in MLLM training and the utilization of hybrid RAG-empowered MLLM agents for data management are discussed in the following subsections.
@@ III-A Cross-Chain Interaction in MLLM Training
In the health center, robust aggregate MLLMs are developed by training on vast amounts of high-quality multi-modal healthcare data[20]. Due to privacy concerns and factors such as patient willingness and incentive structures, hospitals may be reluctant to upload all healthcare data to a central health center[16]. Additionally, the significant computational power required for training MLLMs, combined with the lack of high-performance computing resources at most hospitals, limits the feasibility of using federated learning in this framework[29, 30]. These constraints necessitate alternative approaches for managing data and training models effectively.
In response to these multifaceted challenges, blockchain and cross-chain technologies have emerged as powerful solutions to facilitate secure and decentralized data sharing across healthcare networks. Blockchain technology ensures data integrity and transparency by providing an immutable ledger, while cross-chain technology enables seamless interoperability between diverse blockchain networks, allowing secure data and asset transactions across chains[8]. Recent studies further demonstrate the potential of cross-chain frameworks in significantly enhancing data security and enabling effective collaboration across distributed healthcare systems[16, 8]. To this end, we incorporate cross-chain technology to enable hospitals to securely upload sensitive healthcare data and conduct secure transactions with the health center, ensuring both data privacy and system efficiency.
@@
As shown in Fig. 1, a robust health center utilizes a main chain to manage the comprehensive collection of healthcare data and model updates, and multiple subchains are employed to handle specific tasks from hospitals in diverse regions. These subchains utilize IoMT devices to collect real-time healthcare data from patients, such as temperature, heart rate, and blood pressure, enabling doctors to develop personalized treatment strategies based on a comprehensive analysis of multiple patient attributes[1, 6]. In addition to data collection, the subchains also manage MLLM configurations, and other workflow tasks to support effective healthcare delivery. The main chain ensures centralized oversight, while the sub-chains enable efficient and secure data management operations. Specifically, the main chain $M$ sends data collection tasks to the relay chain $R$. When the subchains $S_{1}$, $S_{2}$, $S_{3}$, and $S_{4}$ receive the tasks, hospitals will upload local multi-modal healthcare data based on the selected contracts (Step a in Fig. 1). Upon successful verification of cross-chain requests by the miners of the relay chain $R$, the relay chain $R$ returns a readiness confirmation, allowing subchains to upload multi-modal healthcare data (Step b in Fig. 1) [16]. After transmitting the multi-modal healthcare data to the main chain $M$, the health center initiates MLLM training (Step c in Fig. 1). Once MLLM training is completed, the health center and hospitals can access the MLLM Application Programming Interface (API) or the weights file of MLLMs through the relay chain $R$ (Step d in Fig. 1), and the health center rewards them with monetary compensation based on their data contribution (Step e in Fig. 1)[16].
@@
Figure 1: The overview of the hybrid RAG-empowered medical MLLM framework for healthcare data management in IoMT. Part A shows the cross-chain interaction for secure healthcare data sharing. Part B depicts the processes of multi-modal input optimization based on a hybrid multi-modal RAG module. Part C presents the framework of MLLM inference based on the multi-modal healthcare data.
III-B Hybrid RAG-empowered MLLM Framework for Data Management in IoMT